OpenStudy (curry):

Diode Problems (Including Zener Diodes) In circuits below with Si diodes (VD0 = 0.7 V) and VZ = 5 V for Zener diodes, A) Find v and/or i for vs = 10 V, B) Find v and/or i for vs ranging from −20 to +20 V.

OpenStudy (curry):

Number 1 has already been completed. @ganeshie8

OpenStudy (curry):

2a) V = 0.7V, I = 0.7/5000 = 0.00014A.

OpenStudy (curry):

2b) Assume off: Vs/10 = i. Vs/2 = V

OpenStudy (curry):

 2b) Assume on: Vs = 5i1 + V, V = Vdo, i1 = id + i